HUBBLE SPACE TELESCOPE DAILY REPORT #2295 PERIOD COVERED: 0000Z (UTC) 01/26/99 - 0000Z (UTC) 01/27/99 Daily Status Report as of 027/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was to be used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. Each iteration of this proposal was affected by the Gyro-3 anomaly that occurred yesterday. (See 2.3) 1.2 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. Each iteration of this proposal was affected by the Gyro-3 anomaly that occurred yesterday. (See 2.3) 1.3 Completed Two Sets of WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C135.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena and they may be used as probes of their environments. They are unique probes of the early Universe. The first iteration of this proposal was affected by the Gyro-3 anomaly that occurred yesterday. (See 2.3) Otherwise, the observations were completed as planned, and no further problems were reported. 1.4 Completed WF/PC-2 6675 (The Evolution of Galaxies in Groups) The WF/PC-2 was to be used to make observations of the radio galaxy PKS0101+02. This proposal was affected by the Gyro-3 anomaly that occurred yesterday. (See 2.3) 1.5 Completed Three Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The first iteration of this proposal was affected by the Gyro-3 anomaly that occurred yesterday. (See 2.3) The successful re-acquisition occurred about five hours into this 8-hour, 40 minute iteration. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.6 Completed Two Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed STIS/CCD/MA1/MA2 7389 (Chromospheres and Transition Layers in Hyades and Pleiades F stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the F class star HD21847 which is located in the Hyades star cluster. This proposal seeks to obtain additional empirical constraints on the heating mechanisms for stellar chromospheres, transition layers and coronae. A wavelength calibration obseration was also taken. The observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Three Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.10 Completed STIS/CCD 7350 (Testing the Supermassive Black Hole Paradigm in Nearby Radio-Quiet AGNs) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the radio-quite active galactic nuclei NGC-3031. Obtaining a large sample of central engine mass determinations is a crucial first step toward understanding the physics of AGN central engines, including the puzzling difference between radio-loud and radio-quiet AGNs. A wavelength calibration observation and a CCD flat were also taken. The observations were completed as planned, and no anomalies were reported. 1.11 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.12 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local group galaxies offer the opportunity for a closer look at young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.13 Completed Two Sets of WF/PC-2 6633 (Cusps or Cores in Barred Bulges?) The WF/PC-2 was used to make observations of the SB class (Barred) galaxies NGC-2950 and NGC-2880in order to study their nuclear bulge region and cusps. These are a highly tri-axial, tumbling, low-mass stellar systems. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. After the Gyro-3 motor current flagged out-of-limits high at 025/2354Z (see yesterday's report for details), the flight software autonomously took Gyro-3 out of the control loop. A gyro bias error of 0.056 deg/min resulted. A 3-gyro HLGBU was uplinked at 026/0155Z, and an FHST map obtained at 026/0240Z initially calculated an attitude error of 5 to 6 degrees. Since the stars could not be positively identified from the map, the coarse attitude (sun sensors/magnetometers) was utilized to reduce the error to 1 to 2 degrees. An ARU/PRT was uplinked at 026/0600Z. Another FHST map taken at 026/0615Z indicated an error of 54 arcsec, too large for a re-acquisition search radius. Accordingly, another ARU/PRT was uplinked at 026/0650Z. The re-acquisition at 026/0655Z was successful and regular science operations commenced. A type T2 slew was performed at 026/1120Z. The high mode bias offset error was good. Subsequent acquisition/re-acquisition activity was nominal. The proposals described in 1.1 through 1.5 were either entirely or partially affected by these events. The HST continues in a 3-gyro configuration. Per an operations request, the NICMOS FPA sensors were enabled at 026/2054Z. As described in HSTAR 6947, batteries 5 and 6 temperatures flagged out-of-limits high at 026/2106Z. A SPA was taken off line to reduce the charge current at 027/0313Z. The temperature parameters returned to normal ranges. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH